OpenStudy (anonymous):

FInd all values of x 3sinx=1+cos2x

OpenStudy (anonymous):

3*2*(sinx(x/2) )cos(x/2) - 2(cos(x/2))^2 = 0

OpenStudy (anonymous):

2cos(x/2) *(3sin(x/2) - cos(x/2)) =0 thus cos(x/2)=0 means x/2 =(2n-1)(pi/2 or x= (2n-1)(pi)

OpenStudy (anonymous):

can you expain how you got the first equation from the beginning equation?

OpenStudy (anonymous):

sinx =2*sin(x/2)*cos(x/2) and cosx= 2(cos(x/2))^2 -1

OpenStudy (anonymous):

other values can be found from (3sin(x/2) - cos(x/2)) =0

OpenStudy (anonymous):

@matricked im still unsure :/ im getting confused by the way its written @robtobey can you help explain? i kinda still dont unerstand

OpenStudy (science0229):

There is a trigonometric identity, called double angle identity. It's\[\sin(2x)=2\sin x \cos x\]\[\cos(2x)=\cos^2x-\sin^2x=2\cos^2x-1=1-2\sin^2x\]

OpenStudy (science0229):

Also, for tangent\[\tan(2x)=\frac{ 2\tan x }{ 1-\tan^2x }\]

OpenStudy (science0229):

Because there is sin(x) at the left side and cos(2x) on the right side, it would be the easiest to use the identity cos(2x)=1-2sin^2x to put everything in terms of sin(x)

OpenStudy (anonymous):

Refer to the Mathematica attachment.

OpenStudy (science0229):

So the equation becomes \[3\sin x=1+1-2\sin^2x=2-2\sin^2x\]

OpenStudy (science0229):

\[2\sin^2x+3\sin x-2=0\rightarrow (\sin x+2)(2\sin x-1)=0\]

OpenStudy (science0229):

Can you take it from here?

OpenStudy (science0229):

@El3th1897

OpenStudy (anonymous):

thank you for replying @science0229 and yea i make the 2 factored equations equal zero then find sinx on both to find the degrees of sinx right?

OpenStudy (science0229):

yeah.

OpenStudy (anonymous):

the first one it becomes sin(x)=-2 there is no solution for that right?

OpenStudy (anonymous):

@science0229

OpenStudy (science0229):

yeah

OpenStudy (anonymous):

cool

OpenStudy (science0229):

you just have to find the solution for sin(x)=1/2.

OpenStudy (science0229):

I'm sure that you can do it, right?

OpenStudy (anonymous):

\[\Pi/6 and 5\Pi/6 \right?\]

OpenStudy (anonymous):

thats in radians pi/6 and 5pi/6

OpenStudy (science0229):

over the interval of [0,2pi].

OpenStudy (science0229):

have to eat. sorry. different time zone here( seoul time zone) fyi. the general solution is (pi/6+2npi) and (5pi/6+2npi) where n=integer